OpenStudy (anonymous):

Prove that the definite integral [sqrt(x^4+1)]dx , 1

OpenStudy (anonymous):

\[\int\limits_{1}^{3}\sqrt{x^{4}+1} dx \ge \frac{ 26 }{ 3 }\]

OpenStudy (tkhunny):

First, try a few things. \(\int\limits_{1}^{3}x^{4}+ 1\;dx = \dfrac{252}{5}\) Okay, that's REALLY large. \(\int\limits_{1}^{3}\sqrt{2}\;dx = 2\sqrt{2}\) Okay, that's prety small. \(\int\limits_{1}^{3}\dfrac{\sqrt{82} - \sqrt{2}}{2}(x-1) + \sqrt{2}\;dx = \sqrt{2}+\sqrt{82}\) That's a little large. I think we might need to break it up into convenient chunks.

OpenStudy (anonymous):

like a cake!

OpenStudy (tkhunny):

Well, maybe, but don't try circular wedges unless you are particularly hardy.

OpenStudy (anonymous):

all I know is i've tried to apply numerous properties but have gotten nowhere

OpenStudy (tkhunny):

Did you try left-hand rectangles at x = 1 and x = 2?

OpenStudy (anonymous):

unfortunately, the problem calls for usage of only properties, not useful methods like Riemann sums

OpenStudy (tkhunny):

Fair enough. Kind of hair-splitting, but let's see where it goes. Please name the "Properties" we get to use.

OpenStudy (anonymous):

essentially these: http://www.emathhelp.net/notes?nid=481

OpenStudy (tkhunny):

Okay, two things come immediately to mind: 1) The "Adjacency of Integral" property essentially defines the fundamentals of finite Reimann Sums, so I still like my first answer. 2) Comparison Property 5, with Example #3 is almost exactly your original problem. Why not try that?

OpenStudy (anonymous):

Comparison property 5 was my first thought as well, and I couldn't get an answer large enough to prove it's >= 26/3. And using the adjacency property would require me to break it up into LOTS of parts because the real value of the integral is 8.98, only 0.31 large than 26/3

OpenStudy (tkhunny):

Okay, then how about Comparison Property #3 \(\sqrt{x^{4} + 1} > x^{2}\) \(\lim\limits_{x\rightarrow\infty}\sqrt{x^{4}+1} = x^{2}\) The key to this solution was the '3' in the denominator. It seemed suspicious or arbitrary.